Fred begins walking toward John's house at 3mi/hr. John leaves his house at the same time and walks toward Fred's house on the same path at a rate of 2mi/hr. How long will it before they meet if the distance between the houses is 4 miles?
Answer by ptaylor(1894)

Let t=time spent walking until they meet
distance(d)=rate(r) times time(t) or d=rt;r=d/t and t=d/r
distance John walked=rt=3*t
distance Fred walked =rt=2*t
Now we know that when they meet, the distance John had walked plus the distance Fred had walked must equal 4 mi. So our equation to solve is:
3t+2t=4
5t=4 divide both sides by 5
t=4/5 hr or 48 min-------------time elapsed until they meet

CK
3*(4/5)+2*(4/5)=4 mi
12/5+8/5=4
20/5=4
4=4

Hope this helps---ptaylor